### v1

# A First Look at Information Highlighting in Stack Overflow Answers

Abstract
Context: Navigating the knowledge of Stack Overflow (SO) remains challenging. To make the posts vivid to users, SO allows users to write and edit posts with Markdown or HTML so that users can leverage various formatting styles (e.g., bold, italic, and code) to highlight the important information. Nonetheless, there have been limited studies on the highlighted information.
Objective: We carried out the first large-scale exploratory study on the information highlighted in SO answers in our recent study. To extend our previous study, we develop approaches to automatically recommend highlighted content with formatting styles using neural network architectures initially designed for the Named Entity Recognition task.
Method: In this paper, we studied 31,169,429 answers of Stack Overflow. For training recommendation models, we choose CNN and BERT models for each type of formatting (i.e., Bold, Italic, Code, and Heading) using the information highlighting dataset we collected from SO answers.
Results: Our models based on CNN architecture achieve precision ranging from 0.71 to 0.82. The trained model for automatic code content highlighting achieves a recall of 0.73 and an F1 score of 0.71, outperforming the trained models for other formatting styles. The BERT models have even lower recalls and F1 scores than the CNN models. Our analysis of failure cases indicates that the majority of the failure cases are missing identification (i.e., the model misses the content that is supposed to be highlighted) due to the models tend to learn the frequently highlighted words while struggling to learn less frequent words.
Conclusion: Our findings suggest that it is possible to develop recommendation models for highlighting information for answers with different formatting styles on Stack Overflow.
keywords:
Stack Overflow, Information highlighting, Named entity recognition, Deep learning
[…]
[1]organization=Department of Computer Science, University of Manitoba, country=Canada
[…]
[2]organization=School of Computing, Queen’s University, country=Canada
[…]
[3]organization=Department of Computer Science and Software Engineering, Concordia University, country=Canada
[…]
[4]organization=Huawei, Canada, country=Canada
1 Introduction
Technical question and answer (Q&A) sites such as Stack Overflow (SO) have become increasingly important for software developers to share knowledge and contribute to communities. Despite Stack Overflow’s success and prevalence, navigating the knowledge on it remains challenging Gottipati et al. (2011); Nadi and Treude (2020); Xu et al. (2017); Zhang et al. (2021a). The previous study shows that finding answers in long posts remains one of the challenges Nadi and Treude (2020); Xu et al. (2017). 37% of all questions on Stack Overflow have more than one answer, and the average length of an answer is 789 characters Nadi and Treude (2020). To make the posts vivid to users, Stack Overflow platform allows users to edit their posts with Markdown and HTML Overflow (2022); StackExchange (2023), so that users can leverage various formatting styles (e.g., Bold, Italic, and Code) to highlight text and direct other users’ attention toward the most important information within posts.
Information highlighting has demonstrated its effectiveness in various domains Nguyen et al. (2015); Ramírez et al. (2019); Strobelt et al. (2015); Wilson et al. (2016); Wu and Yuan (2003), including in Software Engineering Sarkar (2015) (e.g., it saves the reading time of humans and helps programming novices with code comprehension with syntax highlighting). However, little is known about how information is highlighted on technical Q&A sites (e.g., Stack Overflow). For example, how prevalent is the information highlighted? What content is highlighted using different formatting styles? Understanding this could provide a landscape of the usage of information highlighting on technical Q&A sites and shed light on the information that is considered important to developers.
Previous study: In our previous study Ahmed et al. (2022), we performed the first large-scale study on the five most commonly used information highlighting types, which are Bold, Italic, Code, Delete, and Heading, to understand their characteristics and what information is highlighted in the text description of SO answers with them [^1] Note that an SO answer may contain both text and code block, we focus on the information highlighting in text description.. Our analysis of 55,209,643 information highlighting instances across 14,845,929 answers on Stack Overflow reveals that: Overall, information highlighting is prevalent on SO, i.e., 47.6% (14,845,929 out of 31,169,429) of the answers use the studied formatting types to highlight information. 38.5%, 11.3%, and 7.2% of the answers use Code, Bold, and Italic, respectively, and their highlighted content is short (median length is one word). Code formatting is mainly used to highlight source code content, such as identifiers (63.5%). Code is also used to highlight content other than source code, such as Software (4.9%) and Equation (5.2%). Italic and Bold are frequently used to highlight source code content, as well as content that warns about the context where the provided solution works or does not work, updates on answers, and references to an internal or external resource.
Extention: Since information highlighting is prevalent on SO, however, identifying the appropriate content to highlight can be a challenging task for SO users (typically for new users), as they must properly pinpoint the focal points and select an appropriate formatting style, which can be time-consuming and require expertise. Recommending important information to highlight in a post can help improve the presentation of the post and make the important information more visible to SO users. Therefore, in this paper, we extend our recent study Ahmed et al. (2022) by investigating the potential of recommending highlighted content automatically. To do so, we adapt two models (i.e., CNN and BERT) to automatically identify highlighted content for each type of formatting. More specially, we train NER models on each formatting (i.e., Heading, Bold, Italic, and Code) type and recommend the content to highlight. CNN models outperform Bert-based models. CNN models achieve precision values ranging from 0.71 to 0.82. The CNN models for automatic source code content highlighting achieve a recall score of 0.73 and an F1 score of 0.75, outperforming the models for other formatting types. By analyzing the failure cases, we observed that the majority of the failure cases are due to missing identification (i.e., the model misses the content that is supposed to be highlighted). Our analysis shows that the models tend to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge).
Paper Organization. Section 2 presents the background and the related work. Section 3 explains our research questions and how we prepare our data to answer research questions. In Section 4, we present the results of our research questions. Section 5 discuss the implications of our study and threats to validity. Finally, Section 6 concludes our study.
2 Background & Related work
2.1 Background
Stack Overflow allows users to use Markdown and HTML to write and edit posts Overflow (2022); StackExchange (2023). Certain formatting types are used to highlight the information in text description with Markdown and HTML tags. In this study, we focus on the five most commonly used formatting types, which are Bold, Italic, Delete, Code, and Heading. In Table 1, we present the HTML tags and their equivalent Markdown syntax for each formatting type. We group HTML tags and Markdown formatting based on their rendering effect. For example, we group <b> and <strong> as Bold since they render the same effect in the browser when users read the posts on Stack Overflow. In this study, we consider a highlighted sequence of words as an highlighted instance. One post could have multiple highlighted instances. For instance, in Figure 1, the post has 10 Code instances [^2] https://stackoverflow.com/questions/32402475.
Figure 1: An example of SO answers in which the code-related content is highlighted.
In the rest of this paper, we refer to the formatting types Bold, Italic, Heading, and Delete as Text formatting for simplicity’s sake. We use Code and Code formatting, Text and Text formatting exchangeably.
Table 1: The studied formatting types that are used to highlight information, their corresponding HTML tags, and equivalent Markdown. HTML tags are grouped based on their rendering effects.
Type HTML tags Equivalent Markdown
Code
<code>
’example’
Bold
<b>, <strong>
**example**, __example__
Italic
<i>, <em>
*example*
Delete
<del>, <s>
None
Heading
<h1>, <h2>, <h3>, <h4>, <h5>, <h6>
# example, ## example, ### example, #### example, ##### example, ###### example
2.2 Related work
2.2.1 Information highlighting
Previous research has explored the benefits of highlighting information in various domains over the last decades Nguyen et al. (2015); Ramírez et al. (2019); Strobelt et al. (2015); Wilson et al. (2016); Wu and Yuan (2003). Wu and Yuan have shown that highlighting could reduce the cognitive load and thus reduce reading time Wu and Yuan (2003). Jorge et al. investigated the impact of highlighting text in the text classification tasks in the machine learning area and found that highlighting is effective in reducing classification effort for humans Ramírez et al. (2019). Similarly, Nguyen et al. developed a tool to explain the output ML models by highlighting the portion of text and demonstrated its effectiveness in model explanation Nguyen et al. (2015). Various studies have been done to investigate the impact of information highlighting in software engineering, such as code comprehension Beelders and du Plessis (2016); Hannebauer et al. (2018); Palma et al. (2022); Sarkar (2015), source code evolution Escobar et al. (2022). For instance, Advait Sarkar investigated the effect of syntax highlighting on program comprehension and its interaction with programming experience Sarkar (2015), and found that syntax highlighting significantly improves task completion time. Different from prior studies that focus on understanding the influence of information highlighting, we focus on understanding the practice of information highlighting on SO.
2.2.2 Knowledge retrieval on Stack Overflow
Several studies have investigated developers’ challenges and solutions in retrieving knowledge from technical Q&A sites and provided some tools to facilitate the process Gottipati et al. (2011); Nadi and Treude (2020); Xu et al. (2017); Zhang et al. (2021a). For instance, Gottipati et al. found that in software forums it is a painstaking process for users to manually search through answers in various threads of posts Gottipati et al. (2011), and they developed a customized search engine to find relevant answers from various software forums. Similarly, Xu et al. developed a technique called AnswerBot, which takes input as the technical question and generates an answer summary for the question Xu et al. (2017). Zhang et al. pointed out an issue of current comment ranking and display mechanism on SO (e.g., a large portion of useful comments are not visible to users) and proposed an alternative ranking mechanism to alleviate the issue Zhang et al. (2021a). Sarah and Treude investigated the possibility of using two existing techniques, wordpartten and lexrank, to identify the essential sentences from a long SO answer Nadi and Treude (2020). Different from prior studies that focus on retrieving relevant or important answers from SO, we investigate what and how information is highlighted on SO. Our study could provide insights to enhance existing techniques.
2.2.3 Recommendation system for Stack Overflow
Previous studies developed different recommendation tools to ease the use of the Q&A sites, such as recommending hyperlinks Li et al. (2016, 2018b), tags He et al. (2022); Maity et al. (2019); Rekha et al. (2014); Wang et al. (2019a, 2018), and similar questions Wang et al. (2019b); Yin et al. (2019). Li et al. developed LinkLiv to recommend hyperlinks Li et al. (2018b), which uses multiple features, including hyperlink co-occurrences in Q&A discussions, and locations (e.g., question, answer, or comment). Wang et al. developed EnTagRec, which combines two complementary lines in the statistics community Bayesian and frequentist Wang et al. (2018). Wang et al. developed SOTagRec combining a convolutional neural network and collaborative filtering model to infer tags for new postings by studying the historical postings and their tags Wang et al. (2019a). Maity et al. developed DeepTagRec Maity et al. (2019), which learns from the content representation of question title and body, and recommends appropriate question tags on Stack Overflow. Wang et al. developed an approach that combines topical interest, topical expertise, and activeness to recommend answers for new questions Wang et al. (2019b). Similar to these studies, we investigate the potential of developing models for recommending information to highlight in SO answers. We adapted Named Entity Recognition (NER) models to identify the content that needs to be highlighted.
3 Research Questions & Data preparation
3.1 Research questions
This study aims to understand how information highlighting is used in SO answers and what content is highlighted using different formatting types. Hence, we formulate our study by answering the following three research questions:
  • 1.
    RQ1: How prevalent is the information highlighting in SO answers?
  • 2.
    RQ2: What types of information are highlighted with Code formatting in SO answers?
  • 3.
    RQ3: What types of information are highlighted with Bold, Italic, Delete, and Heading formatting in SO answers?
  • 4.
    RQ4: Can we recommend the highlighted content in SO answers automatically?
In RQ1, we aim to understand the prevalence of the usage of the studied highlighting formatting and their characteristics. Understanding this can provide practitioners an overview of the landscape of information highlighting in SO answers. We study Code formatting and the rest formatting types (i.e., Bold, Italic, Heading, and Delete) in separated RQs (RQ2 and RQ3) due to their different nature. We study the information highlighted by different formatting types. By understanding this, we can provide insights into what information is important from users’ perspectives and provide insights for future research. For instance, we can provide insight for the downstream research that leverages SO information to facilitate software engineering tasks (e.g., API documentation enrichmentLi et al. (2018a); Treude and Robillard (2016)). Identifying the appropriate content to highlight can be a challenging task for SO users, as they must properly pinpoint the focal points and select an appropriate formatting style, which can be time-consuming and require expertise. For instance, we observe that, in many SO answer posts, certain contents were not initially highlighted in the first version of the post but were highlighted in subsequent versions, such as the second or third. For instance, in the second and third revisions of a post, a user highlighted the code-related content (e.g., “Integer”, “Exception”) to improve the presentation of the post [^3] https://stackoverflow.com/posts/218510/revisions?page=2. Therefore, in RQ4, we explore the feasibility of utilizing machine learning models to recommend which information to highlight.
3.2 Data preparation
To answer the four RQs, we downloaded a data dump of Stack Overflow from the Stack Exchange data dump dated March 2021444https://archive.org/details/stackexchange. The data dump contains details information about posts (i.e., questions and answers), as well as their revision history. In this study, we include all the answers and we ended up with 52,166,061 answer posts.
As the dataset was large for processing, we imported it into MySQL. For each answer post, we first extract the textual content from its body and exclude code block(s). Note that we extracted code block(s) using tag “<pre>”. Next, we apply the regular expressions defined in Table 1 to identify information highlighting in each answer post, using the Python library Regex. In the end, we extract the content that is highlighted by the studied formatting types. We make our replication package public at https://github.com/shaoweiwang2010/REP_2022_Information_highlight_SO.
4 Results
4.1 RQ1: How prevalent is the information highlighting in SO answers?
Approach: To understand the prevalence of information highlighting in SO answers, we compute the percentage of answers that have the studied formatting types (see Table 1) and the percentage of words highlighted in each answer with each formatting. In addition, we compute the basic statistics of information highlighting instances to understand the characteristics of each formatting type. More specifically, we calculate the distribution of instances for each formatting and the number of words highlighted with them.
Results: Overall, 47.6% (14,845,929 out of 31,169,429) of the studied answers have information highlighted. Among all the answers having information highlighted, an average of 10.6% and a median of 7.1% of the text (in words) is highlighted and each answer has an average of 3.9 and a median of 2 highlighting instances.
Table 2: The answer post-wise and highlighted instance-wise statistics of different formatting types.
Type %Answers #Highlighted instance per answer (mean/median/max) %Highlighted words per answer (mean/median/max) %Instances #Words per instance (mean/median/max)
Bold
11.3%
2.0/1.0/241
9.0%/4.1%/100%
12.0%
2.9/1.0/436
Italic
7.2%
1.9/1.0/354
5.9%/2.3%/100%
7.3%
2.9/1.0/477
Delete
0.07%
1.2/1.0/36
18.1%/10.6%/100%
0.04%
15.7/8.0/701
Code
38.5%
3.7/2.0/490
8.7%/6.3%/100%
78.9%
1.4/1.0/4,669
Heading
1.6%
2.1/2.0/168
9.3%/3.9%/100%
1.7%
3.4/2.0/327
Code is used the most frequently among all studied formatting followed by Bold and Italic. Table 2 presents the statistics of different formatting types. We observe that Code is the most frequently used type. 38.5% of the studied answers have content highlighted with Code. Moreover, 78.9% of the highlighted instances are Code. In addition, the studied answers have two Code instances on the median, which is larger than other formatting types except Heading. This finding is expected since users usually discuss programming problems on SO and it is common to highlight source code in the text. Bold and Italic are the most frequently used formats besides Code. They are used in 11.3% and 7.2% of answers, respectively. Delete is rarely used, only 0.07% of the answers use it.
In general, the length of highlighted content is short. The median length of the content highlighted with Code, Bold, Italic, Deleting, and Heading are 1, 1, 1, 8, and 2 words, respectively. Users tend to highlight single word or phrase using Code, Bold, and Italic. Compared with other formatting types, Delete instances are longer.
Information highlighting is prevalent on SO, i.e., 47.6% of the answers use the studied formatting to highlight information. 38.5% of the answers use Code, which is the most frequently used format, followed by Bold (11.3%) and Italic (7.2%). In general, the length of highlighted content is short.
4.2 RQ2: What types of information are highlighted with Code formatting in SO answers?
Approach: In this RQ, we aim to understand what content is highlighted with the Code. First, we randomly sampled 385 Code instances with a 5% interval and 95% confidence level. Since there is no existing terminologies we could reuse for this purpose, we manually performed a lightweight open coding-like process Seaman (1999); Zhang et al. (2019) to identify the type of content highlighted with Code. The process involves three phases and is performed by the first authors (A1 and A2) of the paper. In phase I, A1 and A2 derived a draft list of types based on 50 Code instances. During the phase, the types were revised and refined. In phase II, A1 and A2 independently applied the derived types to the rest 335 samples. They took notes regarding the diffidence or ambiguity during the labeling. During this phase, no new types were introduced. In phase III, A1 and A2 discussed the results from Phase II to resolve any disagreements until a consensus was reached. The coding process has a Cohen’s kappa of 0.8 (measured before starting Phase III), which indicates a substantial level of the inter-rater agreement Viera et al. (2005). Table 3 presents the definition and the corresponding example of the derived types for Code.
Results: Code is mainly used to highlight source code elements, such as identifiers (63.5%), programming language keywords (9.9%), and statements (7.0%). Table 3 presents the derived types of content that is highlighted with Code and the distribution of each type. Code is used the most frequently to highlight identifiers in source code, such as the name of classes, methods, parameters, and variables. We also observe that 59% of the highlighted identifiers appearing in the code block of their corresponding answers. That is said, Code is commonly used to highlight identifiers in text for referring to the corresponding identifiers in the code blocks. This is reasonable since users usually need to refer to a unit in the source code when discussing the solution. In addition, 9.9% and 7.0% of the studied instances are for Keyword, and Statement, respectively.
Code is also used to highlight content other than source code, such as Software (4.9%), Terminology (1.8%), Equation (5.2%), and Version (0.5%). From Table 3, we observe that users also use Code to highlight content other than code. For instance, in 5.2% of the cases, Code is used to highlight content related to equations. For example, in an answer, answerer discuss the time complexity of binary search “This is O(log n)”, which is highlighted using Code [^5] https://stackoverflow.com/questions/17117375. Code sometimes is used to emphasize the names of a software, a framework, and a lib. For instance, the answerer of an answer mentioned “You are using Mysql as DB …” and show a piece of code to demonstrate the database connection between Java and Mysql [^6] https://stackoverflow.com/questions/24586043.
Table 3: The definition and distribution of types of content highlighted with Code formatting.
Type Definition (example) Count (%)
Identifier
The identifier in source code, e.g, “ArrayList”.
245 (63.6%)
Keyword
The keywords in programming languages, e.g., “for” and “public”.
38 (9.9%)
Statement
A statement of source code, e.g., “plot = last.plot()”.
27 (7.0%)
Equation
Mathematical equation, operator, or number, e.g., “O(log n)”.
21 (5.2%)
Software
The name of softwares, frameworks, tools, and libs, e.g., “Tensorflow” and “MySQL”.
19 (4.9%)
Path
Path and files name, e.g., “src/main/java”.
17 (4.4%)
Terminology
Terminologies related to programming, e.g., “hash table”.
7 (1.8%)
Cmd
Command, e.g., “cloud-init init”
5 (1.3%)
Version
Version information of a software, e.g., “62.1”.
2 (0.5%)
Other
Other than the above defined types.
4 (1%)
Although Code is mainly used to highlighted the content related to source code, it is also used to highlight content other than source code, such as Software (4.9%), Terminology (1.8%), Equation (5.2%), and Version (0.5%).
4.3 RQ3: What types of information are highlighted with Bold, Italic, Delete, and Heading formatting in SO answers?
Approach: In this RQ, we aim to understand what content is highlighted with the studied Text formatting (i.e., Bold, Italic, Delete, and Heading). Similar to RQ2, we performed an manual study. We randomly sampled 385 Text instances with a 5% interval and 95% confidence level. We ended up with 177 Bold, 169 Italic, 3 Delete, and 36 Heading instances. We then identified the types of highlighted content by following the methodology used in RQ2. The coding process has a Cohen’s kappa of 0.78 (measured before starting Phase III), which indicates a substantial level of the inter-rater agreement Viera et al. (2005). Table 4 presents the definition and the corresponding example of the derived types for Text formatting.
Table 4: The definition and distribution of different types of content highlighted with Bold, Italic, Heading, Delete.
Type Definition (example) Bold Italic Delete Heading
Update
The update and new edit in the post, e.g., “Update: According to your comments, I think you will want to try gzfile() in read.table()”.
22 (12.4%) 2 (1.2%) 0 0
Equation
Same as the definition of Equation in Table 3.
3 (1.7%) 2 (1.2%) 0 0
Source code
The union of Identifier/Keyword/Statement/Operator/Cmd/Path in Table 3, e.g., “I’d suggest the usage of startupOrder to configure the startup of route though”.
51 (28.8%) 52 (30.8%) 0 1 (2.8%)
Caveat
A reminder or warn of in which context or condition the provided solution works or does not work, e.g., “It won’t work in that shell though, you need to open a new one”.
33 (18.6%) 43 (25.4%) 0 1 (2.8%)
Reference
Reference to internal or external resource, e.g., “<a href="http://api.jquery.com/not/">.not()</a>”.
12 (6.8%) 17 (10.1%) 0 0
Results
Results or output.
3 (1.7%) 2 (1.2%) 0 0
Terminology
Same as the definition of Terminology in Table 3.
12 (6.8%) 21 (12.4%) 0 0
Heading
The title of a section, e.g., “WORKING EXAMPLE”.
27 (15.3%) 2 (1.2%) 0 33 (91.6%)
Options
Options in software, e.g., “you have to enable Less Secure Sign-In in your google account”.
4 (2.3%) 1 (0.6%) 0 0
Extent
A word/phrase to express extent, e.g., “at least with Hibernate it assumes you want to use a global "hibernate" sequence for all tables, which is just stupid.”
4 (2.3%) 14 (8.3%) 0 0
Version
Same as the definition of Version in Table 3.
0 3 (1.8%) 0 0
Delete
Deleted outdated/wrong description, e.g., “Just use KVO KVC”.
0 0 3 (100%) 0
Other
Other than the above defined types.
6 (3.4%) 10 (5.9%) 0 1
Results: Both Bold and Italic formatting are most frequently used to highlight content related to source code. Table 4 presents the distribution of each Text type. We observe for certain types, Bold and Italic share similar patterns. For instance, in 28.8% of the Bold instances and 30.8% of the Italic instances, content related source code (i.e., Source code) is highlighted. That is said, users also frequently use Bold and Italic to highlight source code other than using Code formatting. Users probably use Bold, Italic, and Code exchangeably, although SO suggests users to tag inline code using Code format StackExchange (2023). Interestingly, we also observe in some cases, the community members changed formatting of the code-related content from Bold and Italic to Code. For instance, in an answer, a user change the formatting for a funciton “strncmp()” from Italic to Code [^7] https://stackoverflow.com/posts/18437465/revisions.
Users frequently use both Bold and Italic to highlight Caveat, Reference, and Terminology. In some cases, it is important to mention the condition or context in which a provided solution works (i.e., Caveat). We notice that such information is usually highlighted in answers. For instance, in a SO answer, the answerer reminded readers “Donot forget add jmtp.dll files (that comes up with jmtp download) as a native library for more info see my answer on …” using Bold. [^8] https://stackoverflow.com/questions/6498179. Bold and Italic are used to highlight Reference and Terminology, although Italic is used more often than Bold.
Users tend to highlight the content of types Update and Heading with Bold, while Italic is more likely to be used for the type of Extent. From Table 4, we can see differences between Bold and Italic. For instance, Bold is more frequently used to highlight Update (12.4%) and Heading (15.3%) compare with Italic. Interestingly, users also use Bold to highlight a heading. One possible reason is that the rendering effect for Bold and Heading are similar. Prior studies reveal that answers on SO are easy to become obsolete Ragkhitwetsagul et al. (2019); Zhang et al. (2019). Therefore, it is not surprising to observe that users highlight the update using Text formatting. Different from Bold, Italic is used more frequently to highlight a word/phrase to express the extent (i.e., Extent).
For Delete, it is all used to highlight the obsolete or incorrect content. In terms of Heading, we observe that in 91.6% of the instances, users used it to highlight heading, which is expected. However, we also observe in one case for highlighting Source Code and one case for emphasizing Caveat.
Apart from source code, both Italic and Bold are used frequently to highlight content in types of Caveat, Reference, Terminology, and Update.
4.4 RQ4: Can we recommend the highlighted content in SO answers automatically?
Approach: In the previous RQs, we focused on analyzing the highlighted content on Stack Overflow. As the highlighting is prevalent and the information type can be categorized, some suggestions based on these uses could benefit the user of Stack Overflow. The objective is to construct a recommendation system designed to propose potential words to users, along with suitable tags, with the aim of enhancing the user’s reading experience and comprehension. In this specific RQ, we explore the feasibility of implementing an automated machine learning-based method for content highlighting. This feature would be helpful, particularly for SO users who are inexperienced in the practice of highlighting information and seek to automatically highlight critical content when composing their posts. Specifically, we aim to create a model that can highlight information in an input answer post using appropriate formatting types (Bold, Italic, Delete, Code, and Heading). To achieve this, we use neural network architectures initially designed for Named Entity Recognition (NER) without actually training NER models. Our task shares similarities with NER, such as the short length of labeled (highlighted) tokens and a limited number of text labeling (formatting) options. Moreover, both NER and automated information highlighting can be formed as a sequence labeling task. Note that we train the neural networks from scratch on our information highlighting dataset.
Figure 2: Layers of the NER model
In this RQ, we have used two different structured models for comparison. One model is based on the Convolutional Neural Network (CNN) model provided by spaCySPACY (2023) and another is a transformer model, BERT Devlin et al. (2019).
Figure 2 presents the architecture of our CNN model. In this paper, we use the Token2Vec model SpaCy (2023) provided by a well-known open-source software library for advanced Natural Language Processing (NLP) tasks named spaCy SPACY (2023). We opt to utilize spaCy since it has emerged as the de facto standard for practical NLP due to its speed, robustness, and performance Le Guillarme and Thuiller (2022). The model architecture comprises a token embedding layer (128 dimensions), four convolutional layers, and an output layer. The token embedding layer is named as MultiHashEmbed by spaCy. It features an embedding layer that embeds a number of lexical attributes separately. Then the results are concatenated as a vector semantic representation of a token for passing through a feed-forward network. As it uses hashing to store embeddings in memory, it also supports faster processing of large datasets. The embedding layer is followed by a 4-layered Convolutional Neural Network (CNN), named MaxoutWindowEncoder. Each of the convolutional layers is followed by a max pooling layer. The final layer is a softmax layer for predicting labels on tokens. Training employs cross-entropy loss and the Adam optimizer.
Figure 3: Layers of the NER model by Bert
On the other hand, BERT is a state-of-the-art transformer model pre-trained on a large dataset Devlin et al. (2019). It can be fine-tuned for numerous natural language processing tasks, encompassing question answering, language inference, text classification, and sequence tagging. The architecture of the model is based on a multi-layer bidirectional Transformer encoder developed by researchers of Google. The model considers the contexts of both preceding and subsequent words. The architecture of BERT consists of multiple transformer blocks, each containing a self-attention mechanism for getting the context of the input sequence. In this work, we fine-tuned BERT for our prediction task. We used the pre-trained model from HuggingFace platform Face (2023).
As our HTML highlighting tags were categorized into five distinct types (see Section 2.1 for details), we initially attempted to train a universal model that would accommodate all five types. However, this approach yielded unsatisfactory results, as the model tended to highlight words as code by default while ignoring other formatting types. Consequently, we trained separate models for each highlighting type. Our findings in RQ1 showed that the Delete type was scarcely used in comparison to the other types, leading us to omit it from our analysis in RQ4. Therefore, we train four models, i.e., $Model_{Bold}$, $Model_{Italic}$, $Model_{Heading}$, and $Model_{Code}$ with each of CNN model and BERT models.
Experiment Setup: We construct the dataset for this RQ by selecting posts that contained highlighted words from Section 3. Then we divided the post into sentences and extracted the corresponding formatting tags described in Section 3. In other words, each data point is a highlighted instance (a sentence containing at least one type of information highlighting). When preparing the training data, we consider the content highlighted with different formatting types as the entities in the NER model. We annotate the tagged words with their formatting tags, considering their position in the phrase. For example, in the sample highlighted sentence shown in Figure 2, i.e., “You will have to use client_wait_for().”, “client_wait_for()” is formatted with a code tag. To label this sentence, we assign the label “B-code” to the highlighted word, indicating that it is the beginning of a highlighted Code content, while the other words in the sentence are labeled as O (Outside). If multiple words are formatted, we used I (Interior) and E (Ending) to label the highlighted content’s internal words and the ending word. For example, “you need to run.git push --force --all first” contains more than one word in the highlighted Code content. So the labels for the phrase would be (B-code: “git”), (I-code: “push”), (I-code: “--force”) and (E-code: “--all”). Note that we train separate models for each highlighting type. When preparing the dataset for a particular highlighting type, if a sentence has other highlighting types rather than the one we target, we omit other types of highlighting types from the sentence. Table 5 presents the basic statistics of our dataset for different formatting types.
Table 5: Number of sentences in train and test datasets.
Type Training set Testing set
Bold
2,590,193
646,486
Italic
1,669,344
417,409
Code
9,292,312
2,324,836
Heading
406,946
102,133
Following prior studies, we partitioned the dataset into an 80:20 ratio for training and testing purposes. We trained our models in 3 epochs where the learning rate is 0.001. We employ two training settings.
We use 32 as the size of the batch in the fixed-size setting (i.e., fixed). Models based on BERT architecture are trained in 0.05 learning rate using 16 batches(fixed) in 6 epochs.
Table 6: Models parameters.
Parameter CNN BERT
Epochs
3
6
Learning rate
0.001
0.05
Batch size
32
16
We ran our experiments on a Linux server with four Nvidia RTX 3090 GPUs, AMD Ryzen 48-Core CPU with 256 GB Ram. For simplicity, we call the models that are trained using the incremental size of batches incremental models and those using the fixed size of batches fixed models.
Evaluation: We use precision, recall, and F1-score to measure the performance of models, following prior studies evaluating NER models Esuli and Sebastiani (2010); Jiang et al. (2016). Instead of using the exact match (i.e., a true positive is counted when all of the tokens of a named entity are labeled correctly, both the length and type), we decide to use a partial match Nadeau and Sekine (2007); Tanabe et al. (2005), in which we consider if each individual token is predicted correctly. The partial match allows us to measure the performance of the segmentation aspects of highlighted content when the content has multiple words. More specifically, we calculate the precision and recall as follows:
$Precision=\frac{Number\ of\ words\ that\ are\ correctly\ predicted}{Number\ of\ words\ that\ are\ predicted\ by\ the\ model}$
$Recall=\frac{Number\ of\ words\ that\ are\ correctly\ predicted}{Number\ of\ words\ that\ are\ supposed\ to\ tag}$
For instance, suppose for a long highlighted sequence of words (e.g., sudo netstat -antp | fgrep LISTEN), while only the words sudo netstat and | fgrep LISTEN are highlighted correctly, the words “-antp |” were missing. In the exact match, such a case is considered incorrectly tagged, while in the partial match, the words sudo netstat and | fgrep LISTEN are considered as correctly predicted.
Results: Out of the four trained models, the Code model obtained the highest F1 score of 0.71 (spaCy). Table 7 presents the results of our trained models in terms of precision, recall, and F1 using CNN and BERT sizes of batches. In general, in terms of F1, we observe that the Code models (F1 score is 0.343 and 0.71 using BERT and CNN models) perform better than other models. This is reasonable, since first for code formatting, we have much more training data than other types. Secondly, code typically has certain special characters (e.g., “.”, and “()” in API “class.function()”) and keywords (e.g., “foreach”), making it easy for the model to identify. Our observation indicates that it is easier to build a model to identify code and highlight them on SO posts than other types of formatting.
On the other hand, BERT archives higher precision(0.699 to 0.92) than spaCy models however the recalls are below 0.05 which is the lowest recall in spaCy models. With this low recall, it ends with the F1 scores lower than the spaCy models. Only the BERT model that was trained for Code performs scores 0.343 in F1 Scoring, which is also lower than the code models of spaCy. So it can be said that spaCy models performed better than BERT models.
Table 7: The performance of models on all formatting types Bold, Italic, Heading, Code with different training settings. F denotes the fixed models and I denotes the incremental models.
Type Precision (CNN) Precision (Bert) Recall (CNN) Recall (Bert) F1 (CNN) F1 (Bert)
$Model_{Bold}$
0.76
0.76
0.11
0.025
0.19
0.049
$Model_{Italic}$
0.82
0.92
0.05
0.003
0.10
0.006
$Model_{Heading}$
0.62
0.84
0.21
0.0093
0.31
0.018
$Model_{Code}$
0.78
0.699
0.65
0.228
0.71
0.343
Our CNN-based models achieve an acceptable precision ranging from 0.62 to 0.82 for different formatting types while obtaining low recall. Overall, except for the Code models, other models have higher precision and lower recall. For models for text formatting, Heading, Bold, and Italic models can achieve at least a precision of 0.62 from the models. While models for text formatting struggle from low recall. The results suggest that as long as the text formatting models recommend the content to highlight, the accuracy is acceptable, while the models miss most of the content that is supposed to be highlighted (low recall).
Table 8: Successful examples of content highlighted by our models (using the incremental setting) along with original text for different types of formatting. Note that since our models are trained for one specific type of formatting, therefore we only show the content that is highlighted by the specific model in the examples.
Type Original text Text highlighted with our models (Incremental).
Bold
NOTE : THIS MIGHT ALREADY SOMETHING THAT YOU HAVE TRIED. This is probably due to the fact that you are using –trusted-host=pypi.python.org.
NOTE : THIS MIGHT ALREADY SOMETHING THAT YOU HAVE TRIED. This is probably due to the fact that you are using –trusted-host=pypi.python.org.
Code
For more info on searching for the elements on DOM such as getElementById,querySelector, please refer here.
For more info on searching for the elements on DOM such as getElementById,querySelector, please refer here.
Italic
Side note: Your code is falling prey to what I call The Horror of Implicit Globals because you never declare filename.
Side note: Your code is falling prey to what I call The Horror of Implicit Globals because you never declare filename.
Heading
OPTION 1: GIT STASH
You may postpone all current changes of your files in the git stash cache. After that the files are equal to former checkout state.
OPTION 1: GIT STASH
You may postpone all current changes of your files in the git stash cache. After that the files are equal to former checkout state.
In Table 8, we present several successful examples, in which the corresponding content is highlighted with correct formatting types. In the Code example, the model identifies all code-related content in the post correctly, such as getElementById and querySelector. For Italic, the model can identify “The Horror of Implicit Globals” in the sentence. For Heading, the model successfully identifies the first heading sentence.
Case Study of Failure Cases: We also investigate the failure cases of the models to provide some insights for future research. The failure cases could be categorized into the following three families:
  • 1.
    Misidentification: Although the content that needs highlight is successfully identified, incorrect formatting types are recommended. In other words, the content is not identified by the correct model.
  • 2.
    False Identification: The model identifies the content that is not supposed to highlight.
  • 3.
    Missing Identification: The model misses the content that is supposed to highlight.
We calculate the distribution of three failure cases over each model using the confusion matrix. For instance, to compute the number of missing identification cases from $Model_{Bold}$, we count the number of false negative instances in the confusion matrix on the testing data generated by $Model_{Bold}$.
Table 10 presents the distribution of those three types of failures across different models. Overall, the majority of the failure cases are missing identification (e.g., 55% for incremental model) and false identification (e.g., 39% for incremental model), while misidentification takes the least portion (e.g., 6.2% for incremental model). All models suffer from high missing identification except Code models, which explains the low recall in Table 7, typically for Bold and Italic models. 89% and 66% of the failures in Bold and Italic CNN models are missing identification. One possible explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge) Kandpal et al. (2022); Mireshghallah et al. (2022). To test this assumption, we count the frequency of the words that are correctly predicted and the words that are not identified by models in the training data. Table 9 presents the mean/median frequency of words that are correctly predicted and those that are not identified by models. On top of it, Figure 4 compares their distributions. We observe that the mean and median frequency of words that are predicted correctly is much larger than those of words that are not identified in all models, except the median value of Code.
Table 9: The mean/median frequency of words that are correctly predicted (correct prediction) and those that are not identified by models (missing identification) in different models.
Type Mean/Median of missing identification Mean/Median of correct prediction
Bold
4.05/1.0
193.05/22.0
Italic
6.11/1.0
674.48/130.0
Heading
2.53/1.0
34.73/5.0
Code
1.35/1.0
5.37/1.0
Figure 4: Comparison of the frequency distribution of words that are not identified (missing identification) and those that are correctly predicted (correct prediction) on different models.
Table 10: The distribution of different failure types among CNN models of different formatting types.
Failure Type Formatting Type
Bold
Code
Italic
Heading
Total
Misidentification
6.8%
6.7%
1.6%
12%
6.2%
Missing Identification
89%
39%
66%
59%
55%
False Identification
4%
54%
32%
28%
39%
Code, Italic, and Heading models tend to have false identification failures, especially the Code models. For instance, in example 7 of Table 11, the Code model identify “setTimeout” while it is not highlighted in the original text. setTimeout is indeed a function name and one possible explanation is that setTimeout is frequently highlighted with Code in the training data, and the model can recognize it. Similarly, in example 4, the Italic model identifies “I currently” after “Disclaimer:” as Italic. Recall example 3 in Table 8, in the original text, “Disclaimer: I work at SerpApi” was italicized. The model may identify the word “Disclaimer:” and tend to italicize the content after it.
We also observe a small portion of misidentification cases. For instance, in example 1 in Table 11, the original text is highlighted with Heading, while our model identifies the content but formats it using bold. One possible explanation is that those types of formatting share common usage patterns to some extent and confuse the models. We actually observed such a phenomenon in Section 4.3. For instance, we observe that both Bold and Italic are used to highlight source code and Caveat. Heading is used to highlight source code.
Table 11: Failure cases of spacy detection of tags for all the types Bold, Italic, Heading, and code.
ID Type Original text
Text highlighted with our models (Incremental).
Failure type
1
Bold
Swift 5.2.x
First of all, you need to declare an "easy to use" typealias for your block:
Swift 5.2.x
First of all, you need to declare an "easy to use" typealias for your block:
Misidentification: Detected heading as bold.
2
Bold
If you want to extend it’s expiration date, IT WILL THROW THE EXCEPTION!
If you want to extend it’s expiration date, IT WILL THROW THE EXCEPTION!
False Identification: detected normal texts as bold.
3
Bold
From now, your frontend application will use access token in the Authorization header for every request.
From now, your frontend application will use access token in the Authorization header for every request.
Missing Identification: missed Bold.
4
Italic
Disclaimer: I currently work for Mapbox.
Disclaimer: I currently work for Mapbox.
False Identification: detected normal texts as Italic.
5
Heading
Close Sublime and Quit It. Open up your desired project that you want to use Sublime with in Unity
Close Sublime and Quit It. Open up your desired project that you want to use Sublime with in Unity
False Identification: detected normal texts as Heading.
6
Heading
Because there are no spaces in the text. Chinese is commonly written without any space characters between words or sentences. The ASS subtitle format, having been originally designed by and for European language speakers, unfortunately only breaks words on actual space characters. Your subtitle file does not contain any space characters, so the lines are never broken.
Because there are no spaces in the text. Chinese is commonly written without any space characters between words or sentences. The ASS subtitle format, having been originally designed by and for European language speakers, unfortunately only breaks words on actual space characters. Your subtitle file does not contain any space characters, so the lines are never broken.
Missing Identification: Missing Heading for “Because there are no spaces in the text”.
7
Code
setTimeout seems like the right guy for the job. To off a click use jQuery’s .off() Method. Also, the .one() method to attach a callback only once :
setTimeout seems like the right guy for the job. To off a click use jQuery’s .off() Method. Also, the .one() method to attach a callback only once :
False Identification: Detected normal text as code.
8
Code
Most likely that the website you are trying to scrape is a dynamic website, meaning that it is Javascript generated code and can’t be scraped with only requests and beautifulsoup .
Most likely that the website you are trying to scrape is a dynamic website, meaning that it is Javascript generated code and can’t be scraped with only requests and beautifulsoup .
Missing Identification: Missed the Code tag.
Our models achieve a precision ranging from 0.62 to 0.82 for different formatting types. It is easier to build a model to recommend Code than other types. Models for text formatting types (i.e., Heading, Bold, and Italic) suffer low recall. Our analysis of failure cases indicates that the majority of the failure cases are due to missing identification. One explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge).
5 Discussion
5.1 Implication of our findings
Future search should consider the highlighted content for the downstream tasks that leverage information from the SO answers. In RQ3, we observe that users tend to highlight important information for a provided answer using Text formatting, such as a reminder or warn of in which context or condition the provided solution works or does not work (i.e., Caveat), an update and new edit in answers (i.e., Update). Such content is informative and important to users when learning and applying knowledge from the answers. Therefore, it is substantial to consider such information for the downstream tasks that leverage information extracted from SO answers, such as answer summarization Ren et al. (2019), and API documentation enrichment Li et al. (2018a); Treude and Robillard (2016).
Our findings provide insights to improve future research in developing approaches for automatic information highlight. In RQ4, we investigate the potential of deep learning NER models in recommending information to highlight on SO posts. Code is easier to identify than other types of formatting. Furthermore, all text-related models (i.e., Italic, Bold, and Heading) suffer from low recall. We reveal one reason is long-tail knowledge. The model tends to memorize frequently highlighted words and miss the recognition of the less-frequent (e.g., tail knowledge). Future research is encouraged to leverage techniques Zhang et al. (2021b), e.g., data augmentation and class rebalancing for tail knowledge.
5.2 Threats to validity
Internal Validity: Our study involved qualitative analysis in RQs. To reduce the bias, each instance was labeled by two of the authors, and discrepancies were discussed until a consensus was reached. We also showed that the level of inter-rater agreement of the qualitative studies is high. In RQ4, we evaluate our models using precision, recall, and F-1 score with partial match. Although, there might be other metrics that could be used for our task., those metrics are commonly used in NER models, and we follow previous studies Esuli and Sebastiani (2010); Jiang et al. (2016). In RQ4, we utilize a NER architecture from spaCy and train it from scratch based on our own dataset. It was compared with BERT, a transformer model for understanding the quality of the performances of separate architectures. We encourage future research to use more advanced pre-trained large language models.
External Validity: One external threat is that it is not clear whether our findings still hold on other Q&A websites. We needed to conduct several qualitative analysis in our RQs; however, it is impossible to manually study all instances. To minimize the bias when conducting our qualitative analysis, we took statistically representative random samples of all relevant revisions, in order to ensure a 95% confidence level and 5% confidence interval for our observations.
6 Conclusion
This paper is the first large-scale study of information highlighting in the text description of SO answers. We find that information highlighting is prevalent, with 47.6% of the 31,169,42 studied answers having text highlighted. We propose a terminology to categorize highlighted text in SO answers and find that source code content (e.g., identifiers and programming keywords) are frequently highlighted using Code and other highlighting formats (i.e., Bold and Italic). Besides code, users also tend to highlight updates (e.g., updates of answers), caveats (i.e., a reminder or warning of in which context or condition the provided solution works or does not work), and references. Further studies could put more effort into investigating how to use the highlighted content for downstream tasks (e.g., answer summarizing) that leverage information from SO answers, and provide tools that can suggest highlighted text for SO users. In addition, we also investigate the potential of recommending the content to be highlighted automatically by adopting named entity recognition (NER) models. Our experimental results highlight that our CNN models consistently achieve precision scores between 0.71 and 0.82 across different formatting types, although they exhibit lower recall rates. On the other hand, while BERT demonstrates superior precision, it struggles with even lower recall. It’s worth noting that spaCy’s Code model performs exceptionally well, with an impressive F1 score of 0.71, outperforming other models in our study. Although BERT usually has better performance, in our case the spaCy model based on CNN architecture worked better. This might be attributed to spaCy’s ability to learn specific patterns within our custom dataset that BERT’s transformer model may have missed. In future research, we plan to enhance the transformer models through fine-tuning and work on improving the recall rates for both models.